You deposit $17,000 each year for 10 years at 7%. Then you earn 9% after that. If you leave the money invested for another 5 years how much will you have in the 15th year?

if you can i need all steps
Thank you
1
Expert's answer
2015-01-09T08:56:41-0500
If you deposit $17,000 each year for 10 years at 7%. Then you earn 9% after that. If you leave the money invested for another 5 years, then you will have in the 15th year the next sum:
FV = PV*(1+r)^n = 17,000*(1.07^1 + 1.07^2 + 1.07^3 + 1.07^4 + 1.07^5 + 1.07^6 + 1.07^7 + 1.07^8 + 1.07^9 + 1.07^10)*1.09^5 = $386,688.80